OpenStudy (anonymous):

find values of a and d such that f(x)=asin3x+d has a range of 3

OpenStudy (anonymous):

from 3 to 7 is 4 units, so you want the amplitude to be 2

OpenStudy (anonymous):

in other words you want \[f(x)=2\sin(3x)+d\]

OpenStudy (anonymous):

without the \(d\) part, \(2\sin(3x)\) would go from \(-2\) to \(2\) but you want it to go from \(3\) to \(7\) so lift it up 3 units by making \[f(x)=2\sin(3x)+3\]

OpenStudy (anonymous):

no, that is wrong sorry

OpenStudy (anonymous):

you want it to go from 3 to 7 so you have to lift it up FIVE units

OpenStudy (anonymous):

\[f(x)=2\sin(3x)+5\] will do the job